The probability that event A will occur is 40%, while the probability that event B will occur is 60%.
What is the probability that event A or event B will occur, if the events are indendent?
a. 20%
b. 24%
c. 76%
d. 100%
Thank you .

Found 2 solutions by stanbon, robertb:
Answer by stanbon(75887)

The probability that event A will occur is 40%, while the probability that event B will occur is 60%.
What is the probability that event A or event B will occur, if the events are indendent?
---
If the events are independent P(A and B) = P(A)*P(B) = 0.4*0.6 = 0.24
===============
Cheers,
Stan H.
=============
a. 20%
b. 24%
c. 76%
d. 100%

Answer by robertb(5830)

P(A or B) = P(A) + P(B) - P(A and B)
= P(A) + P(B) - P(A)*P(B) = 0.40 + 0.60 - 0.40*0.60 = 0.76
c. 76%